OpenStudy (brookemariexx):

What is the average rate of change of y with respect to x over the interval [–3, 5] for the function y = 2x + 2? 16 1/2 2 1/4

OpenStudy (brookemariexx):

5?

OpenStudy (brookemariexx):

i tried solving the equation by plugging in the values f(-3) = 2(-3) + 2 *(5) f(-3) = -6 +10 f(-3) = 4 f(5) = 2(10) +2(1) = f(5) = 22-4 18/-8 but not sure where i went wrong

OpenStudy (mhchen):

Oh I'm sorry, it's [-3,5], my bad. Yeah your method is correct, let me check your work.

OpenStudy (brookemariexx):

thank you, i cant seem to find the answer on the answer choices so i know i did something wrong

OpenStudy (mhchen):

I'm not sure why you multiplied f(-3) = 2(-3) + 2 by 5. I think it should only be f(-3) = 2(-3) + 2 f(-3) = -6 + 2 f(-3) = -4 And for f(5): f(5) = 2(5) + 2 f(5) = 10 + 2 f(5) = 12 When you plug in a function like: \[f(x) = x^2+x+3\] and do like f(3), it becomes \[f(x) = (3)^2+(3)+3\]

OpenStudy (mhchen):

So then the average rate of change is \[\frac{12-(-4)}{5-(-3)}=\frac{16}{8}=2\] Right?

OpenStudy (brookemariexx):

ahh thank you so much i understand now!

OpenStudy (mhchen):

yaaaa